OpenStudy (anonymous):

the expression (secx+tanx)2 is the same as ______

Nnesha (nnesha):

\(\huge\color{green}{{\rm welcome}~\rm to~open~study!!:)}\)

Nnesha (nnesha):

for example \(\huge\color{red}{{\rm (a+b)^2}~}\) \(\huge\color{red}{\rm = (a+b)(a+b)}\) its mean you have same two order pair then you have to multiply yu can use foil method

Nnesha (nnesha):

and remember this \(\huge\color{orange}{{\rm (a+b)^2}~}\) = \(\huge\color{orange}{{\rm a^2+}~\rm2ab+b^2}\)

OpenStudy (anonymous):

i"m sorry I'm a little stuck can you do the first step please?

Nnesha (nnesha):

its oky :) okay first you should expand the bracket you know secx= ??? and tanx= ??

OpenStudy (anonymous):

\[1/cosx?\] and \[sinx/cosx?\]

Nnesha (nnesha):

yesh that's right so  they both have same denominator can you solve bracket ??